OpenStudy (anonymous):

Solve equation or formula for indicated variable. S=5r^2t for t?

OpenStudy (anonymous):

@JakeV8

OpenStudy (anonymous):

Is that "t" in the exponent? Like:\[S = 5r ^{2t}\]

OpenStudy (anonymous):

no the t is not an exponent just the ^2

OpenStudy (campbell_st):

simply divide bout sides by 5r^2 to make t the subject.

OpenStudy (anonymous):

i dont know how to divide both sides by 5r^2

OpenStudy (anonymous):

im skipping that question Name the property of real numbers illustrated by the equation.-4(x + 3) = -4x - 12 Distributive Property RIGHT???

OpenStudy (campbell_st):

just treat it like a number say u so S = ut how do you make t the subject...?

OpenStudy (anonymous):

\[S = 5r ^{2}t\] \[\frac{ S }{ 5r ^{2} } = \frac{ 5r ^{2}t }{ 5r ^{2} }\]

OpenStudy (anonymous):

yep i dont even see where you got u

OpenStudy (anonymous):

ok that makes since jake

OpenStudy (anonymous):

Yes, what @campbell_st said is right too... don't get confused by having several variables. To solve for one variable, you just get all that other stuff out of the way. In this problem, you needed to divide by that "stuff" (the "stuff" was 5r^2) to get it away from the "t" that you were trying to isolate. But you have to do the same thing to both sides, so you also divided S by 5r^2.

OpenStudy (anonymous):

ok then its S5r^2 = t or no

OpenStudy (campbell_st):

nope... not correct.... and re-read the posts and you'll see a solution.

OpenStudy (anonymous):